OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2 . Answer y - 6 = 1/2 (x + 3) y = 1/2 x - 6 y+ 3 =1/2 (x - 6) x - 2y = 12

sam (.sam.):

m=1/2 y=-3 x=6 use \[y-y_1=m(x-x_1)\] \[y-(-3)=1/2(x-(6))\]